# Image Captioning via Compact Bidirectional Architecture

### v2

@@ Abstract
Index Terms:
Image Captioning, Compact Bidirectional Architecture, Transformer, LSTM.I Introduction
@@
In this work, we introduce a Compact Bidirectional Transformer model for image captioning (dubbed as CBTrans) that can leverage bidirectional context implicitly and explicitly while the decoder can be executed parallelly. Specifically, it is implemented by tightly coupling the L2R and the R2L flows into a single compact model and optionally allowing explicit interaction between the two flows (as shown in Fig. 1(d)). The parallelism of our decoder is obvious since it is composed of a single Transformer decoder. We claim our decoder can exploit bidirectional context explicitly since we optionally allow interaction between L2R and R2L flows and implicitly since one copy of shared parameters has the ability of supporting both L2R and R2L decoding. Let’s take Fig. 1(d) as an example. During training, each image is associated with two captions instead of single caption as in the conventional L2R captioning models. One caption is from left to right with a prefix and the other one is from right to left with a prefix. Inside the CBTrans model, the generation of a target word (e.g., ’walking’) can not only depend on previous words in its own flow (e.g.,’a man in a hat is’,i.e., past context) but also optionally previous words in the other flow (e.g.,’beach the on horse a with’, i.e., future context). The joint loss is composed of both L2R and R2L losses and the model is end-to-end trainable. During inference, CBTrans model takes both and as the text input in the first step and optionally allows interaction between the two flows along the entire decoding process. Finally, the output caption of L2R ’flow’ and the one of R2L ’flow’ are ranked based on their probabilities and the larger one is chosen as the output (i.e. sentence-level ensemble).
We conduct extensive ablation studies on the MSCOCO benchmark dataset to better understand and verify the effectiveness of this model. We find that the compact architecture serves as a good regularization for implicitly exploiting bidirectional context and a single CBTrans model achieves an effect of sentence-level ensemble, which usually needs to train and save two models for improving final predictions as shown in Fig. 1(c). By combining with word-level ensemble, the effect of sentence-level ensemble is further enlarged. Overall, the compact bidirectional architecture and sentence-level ensemble play more important roles than the explicit interaction mechanism. We further extend the conventional one-flow self-critical training to the two-flows version under this model architecture and achieve new state-of-the-art results in comparison with non-vision-language-pretraining models. Finally, we verify the generality of this compact bidirectional architecture by extending it to LSTM backbone and proposing the CBLSTM model.
@@ II-B Multi-task Learning
Multi-task learning is a useful learning paradigm to improve the supervision and the generalization performance of a task by jointly training it with related tasks [36]. Here, we only review some multi-task learning works related to captioning. To the best of our knowledge, there are roughly two types so far. One is to jointly train the task of captioning and syntax generation (e.g. Part-of-Speech) [37, 38, 39, 40]. The other one is to train a unified captioning model on multilingual (e.g. English and Chinese ) captioning datasets [41, 42, 43, 44]. Different from the above multi-task learning methods, one task is the L2R generation and the other one is the R2L generation in our method.
@@ III Approach
In this section, we first present two representative compact bidirectional architectures built on the well-known Transformer [15] and LSTM [51] and then introduce the training procedure for model optimization.
@@ IV Experiments
Each image is represented as a set of region features with dimensions extracted by object detector [52]. We try two kinds of features,i.e.,Up-Down feature with regions per image [25],and VinVL feature with up to regions per image [34], which is extracted by a more powerful detector. The dictionary is built by dropping the words that occur less than times and ends up with a vocabulary of . Captions longer than words are truncated. Our CBTrans model almost follows the same model hyper-parameters setting as in [15] (). And the word embedding size, LSTM hidden state size and image feature projection size of our CBLSTM model are all set to . As for the training process, we train CBTrans (CBLSTM) under cross-entropy loss for () epochs with a mini batch size of (). The optimizer in [15] is used with a learning rate initialized by 5e-4 and the warmup step is set to for CBTrans model and Adam [72] optimizer is used with learning rate initialized by 5e-4 and decayed by a factor 0.8 every three epochs for CBLSTM model. We increase the scheduled sampling probability by 0.05 every 5 epochs. We then optimize the CIDEr score with self-critical training for another 15 (20) epochs with an initial learning rate of 1e-5 (5e-5) for CBTrans (CBLSTM) model. During testing, unless stated otherwise, we use the standard beam search for conventional Transformer and Up-Down models with beam size , which is the best width based on our observation. It is worth noting that the beam search method for CBTrans and CBLSTM models is a little different from the standard one. Specifically, if the total beam size is , then both L2R and R2L flows should independently keep a standard beam search with beam size for each step. Unless stated otherwise, the beam size of each flow is set to for CBTrans and CBLSTM models.
@@ IV-C2 What is the effect of each component that constitutes the compact bidirectional architecture?
To further investigate the effectiveness of our models and each component, we conduct substantial ablation studies, as shown in Table IV. Before diving into the details, we emphasize that our model doesn’t increase any model parameters except two special symbols (i.e., and ) embedding compared to the standard baseline model and they share the same training script.
@@ V Conclusion
In this paper, we introduce a compact bidirectional Transformer model for image captioning (CBTrans) that can leverage bidirectional context implicitly and explicitly and the decoder can be executed parallelly. We conduct extensive ablation studies on MSCOCO benchmark and find that the compact architecture, which serves as a regularization for implicitly exploiting bidirectional context, and the sentence-level ensemble play more important roles than the explicit interaction mechanism. We further propose to combine word-level and sentence-level ensemble seamlessly and extend the conventional self-critical training to achieve new state-of-the-art results in comparison with non-vision-language-pretraining models. Finally, we verify the generality of this compact bidirectional architecture by extending it to LSTM backbone. The proposed method is orthogonal to other sophisticated methods for image captioning, including vision-language pretraining, and we leave integrating these methods for better performance as our future work.

### v3

@@ Abstract
I Introduction
@@
In this work, we introduce a Compact Bidirectional Transformer model for image captioning (dubbed CBTrans) that can leverage bidirectional context implicitly and explicitly while the decoder can be executed in parallel. Unlike refinement-based models that rely on two sequentially executed networks, our model is termed compact because it integrates both L2R and R2L decoding flows into a single Transformer or LSTM network with shared parameters. This design not only preserves parameter efficiency but also enables parallel bidirectional decoding within a unified architecture, allowing the decoder to exploit future context without requiring a separate refinement stage. Specifically, it is implemented by tightly coupling the L2R and the R2L flows into a single compact model and optionally allowing explicit interaction between the two flows (as shown in Fig. 1(d)). During training, each image is associated with two captions instead of a single caption as in conventional L2R captioning models. One caption is from left to right with a prefix and the other one is from right to left with a prefix. Inside the CBTrans model, the generation of a target word (e.g., ’walking’) can not only depend on previous words in its own flow (e.g.,’a man in a hat is’,i.e., past context) but also optionally previous words in the other flow (e.g.,’beach the on horse a with’, i.e., future context). The joint loss is composed of both L2R and R2L losses and the model is end-to-end trainable. During inference, CBTrans model takes both and as the text input in the first step and optionally allows interaction between the two flows along the entire decoding process. Finally, the output caption of L2R ’flow’ and the one of R2L ’flow’ are ranked based on their probabilities and the larger one is chosen as the output (i.e. sentence-level ensemble).
We conduct extensive ablation studies on the MSCOCO benchmark dataset to better understand and verify the effectiveness of this model. We find that the explicit interaction slightly contributes to the final performance when a proper hyper-parameter is selected. We find that the compact architecture serves as a good regularization for implicitly exploiting bidirectional context and a single CBTrans model achieves an effect of sentence-level ensemble, which usually needs to train and save two models for improving final predictions as shown in Fig. 1(c). By combining with word-level ensemble, the effect of sentence-level ensemble is further enlarged. Overall, the compact bidirectional architecture and sentence-level ensemble play more important roles than the explicit interaction mechanism. We further extend the conventional one-flow self-critical training to the two-flows version under this model architecture and achieve new state-of-the-art results in comparison with non-vision-language-pretraining models. Finally, we verify the generality of this compact bidirectional architecture by extending it to LSTM backbone and proposing the CBLSTM model.
@@ II-B Multi-task Learning
Multi-task learning is a useful learning paradigm to improve the supervision and the generalization performance of a task by jointly training it with related tasks [36]. Here, we only review some multi-task learning works related to captioning. To the best of our knowledge, there are roughly two types so far. One is to jointly train the task of captioning and syntax generation (e.g. Part-of-Speech) [37, 38, 39, 40]. The other one is to train a unified captioning model on multilingual (e.g. English and Chinese ) captioning datasets [41, 42, 43, 44, 45]. Different from the above multi-task learning methods, one task is the L2R generation and the other one is the R2L generation in our method.
@@ III Approach
In this section, we first present two representative compact bidirectional architectures built on the well-known Transformer [15] and LSTM [52] and then introduce the training procedure for model optimization. The compact nature of our architecture stems from the shared parameters between L2R and R2L flows, allowing bidirectional decoding within a single network, as opposed to dual-network sequential designs. There is no increase in the number of model parameters during this process.
@@ IV Experiments
Each image is represented as a set of region features with dimensions extracted by object detector [53]. We try two kinds of features,i.e.,Up-Down feature with regions per image [25],and VinVL feature with up to regions per image [34], which is extracted by a more powerful detector. The dictionary is built by dropping the words that occur less than times and ends up with a vocabulary of . Captions longer than words are truncated. Our CBTrans model almost follows the same model hyper-parameters setting as in [15] (). And the word embedding size, LSTM hidden state size and image feature projection size of our CBLSTM model are all set to . As for the training process, we train CBTrans (CBLSTM) under cross-entropy loss for () epochs with a mini batch size of ().
The optimizer in [15] is used with a learning rate initialized by 5e-4 and the warmup step is set to for CBTrans model and Adam [73] optimizer is used with learning rate initialized by 5e-4 and decayed by a factor 0.8 every three epochs for CBLSTM model. We increase the scheduled sampling probability by 0.05 every 5 epochs. We then optimize the CIDEr score with self-critical training for another 15 (20) epochs with an initial learning rate of 1e-5 (5e-5) for CBTrans (CBLSTM) model. During testing, unless stated otherwise, we use the standard beam search for conventional Transformer and Up-Down models with beam size , which is the best width based on our observation. It is worth noting that the beam search method for CBTrans and CBLSTM models is a little different from the standard one. Specifically, if the total beam size is , then both L2R and R2L flows should independently keep a standard beam search with beam size for each step. Unless stated otherwise, the beam size of each flow is set to for CBTrans and CBLSTM models.
It is worth noting that during beam search for CBTrans/CBLSTM, the L2R and R2L flows decode in lockstep. At each timestep, both flows generate one token based on their own past and the other flow’s past. If one flow finishes earlier (predicts an token), its sequence remains static for subsequent steps while the other continues, allowing the longer flow to still leverage the completed ‘future context’ from the finished flow via the bidirectional attention mechanism.
@@ IV-C2 What is the effect of each component that constitutes the compact bidirectional architecture?
To further investigate the effectiveness of our models and each component, we conduct substantial ablation studies, as shown in Table IV. Before diving into the details, we emphasize that our model doesn’t increase any model parameters except two special symbols (i.e., and ) embedding compared to the standard baseline and they share the same training script.
@@ V Conclusion
In this paper, we introduce a compact bidirectional Transformer model for image captioning (CBTrans) that can leverage bidirectional context implicitly and explicitly and the decoder can be executed parallelly. We conduct extensive ablation studies on MSCOCO benchmark and find that the compact architecture, which serves as a regularization for implicitly exploiting bidirectional context, and the sentence-level ensemble play more important roles than the explicit interaction mechanism. We further propose to combine word-level and sentence-level ensemble seamlessly and extend the conventional self-critical training to achieve new state-of-the-art results in comparison with non-vision-language-pretraining models. Finally, we verify the generality of this compact bidirectional architecture by extending it to LSTM backbone. The proposed compact bidirectional architecture is orthogonal to vision-language pretraining approaches, which focus on large-scale cross-modal pretraining. Our decoder could be integrated into vision-language pretraining frameworks by replacing their unidirectional decoders with our bidirectional design for further studies, potentially enhancing caption quality through better context utilization while retaining the benefits of pretrained representations.